evaluate each expression.
This is the problem I need help with:
((( (a/b)^2 + b^3 for a= 24 and b= -6)))

Answer by gonzo(654)

formula given is (a/b)^2 + b^3
a = 24
b = -6
a/b = 24/(-6) = (-4)
(-4)^2 = 16
b^3 = (-6)*(-6)*(-6) = 36*(-6) = (-216)
answer is 16 - 216 = -200.
working it through directly we get
(a/b)^2 + b^3
= (24/-6)^2 + (-b)^3
= (-4)^2 + (-6)^3
= 16 - 216
= -200.